OpenStudy (anonymous):

the sum of three consecutive odd numbers is 27 find the middle number

OpenStudy (anonymous):

middle number is the average of the 3, that means 27/3=9

OpenStudy (anonymous):

thank you i wanted to be sure

OpenStudy (anonymous):

could you help me with another one

OpenStudy (anonymous):

yeah, sure

OpenStudy (anonymous):

find the number such that twice it's square is equal to two times it's original number

OpenStudy (anonymous):

2 x^2 =2x or, 2x(x-1)=0 x=0,1